Complete each ordered pair so that it satisfies the given
equation.
Y= 1.8x +22.6: (1, ), (-10, ), ( , 22.6)

Answer by stanbon(75887)

Y= 1.8x +22.6:
(1, )
f(1) = 1.8+22.6 = 24.4
------
(-10, )
f(-10) = 1.8(-10)+22.6 = -18+22.6 = 4.6
------
( , 22.6)
22.6 = 1.8x+22.6
1.8x = 0
x = 0
==================
Cheers,
Stan H.
=======